OpenStudy (steve816):

What are the solutions to this? 3x⁴ + 4x³ - 36x² + 64 = 0

Nnesha (nnesha):

when you're not able to factor by grouping then you should use p/q method \[\rm \frac{ p }{ q }=\frac{ \pm~all~factors ~of~constant~term }{ \pm~all~factors~of~leading ~coefficient }\]

Nnesha (nnesha):

^to find all possible zeros and then we have to test each possible zero by either using synthetic division or replacing x with each number if you get 0 as final answer then the number u used would be the real zero

Nnesha (nnesha):

and yes that's correct

OpenStudy (steve816):

That seems like a tedious task...

Nnesha (nnesha):

why did you delete ur comment? and the arrow poimting out my first comment

Nnesha (nnesha):

n*

OpenStudy (steve816):

because i thought it was wrong sorry

OpenStudy (steve816):

so do I have to plug in all the numbers till I get 0?

Nnesha (nnesha):

yes correct. but first write all factors of 64 and 3

OpenStudy (steve816):

ohh noooo is there a easy way?

Nnesha (nnesha):

you can do synthetic division.

Nnesha (nnesha):

and the other easy way is graphing....

OpenStudy (steve816):

oh okay, I know how to do synthetic division, but which value do I divide by?

Nnesha (nnesha):

that would be the 2nd step first divide factors of constant by factors of leading coefficient

Nnesha (nnesha):

hat are the factors of 64 ?

Nnesha (nnesha):

what**

OpenStudy (steve816):

64 32 16 8 4 2 1

Nnesha (nnesha):

correct and the factors of 3 are 1 and 3 that's it \[\frac{ \pm 1 , \pm 2, \pm 4 ,\pm 8 ,\pm 16,\pm 18 , \pm 32 , \pm 64}{ \pm 1 , \pm 3 }\] we don't know if its negative or positive so we should keep both sign and now divide all numbers at the numerator by each number in the denoinator

Nnesha (nnesha):

denominator ** and cross out duplicates numbers u get after division :)

OpenStudy (steve816):

What do I do once I find all the numbers?

Nnesha (nnesha):

then you have to test each number which number would give you the value of 0 as a final answer ?

Nnesha (nnesha):

there is a typo 18 which isn't a factor it should be \[\frac{ \pm 1 , \pm 2, \pm 4 ,\pm 8 ,\pm 16,\pm 32 , \pm 64}{ \pm 1 , \pm 3 }\] divide all the top numbers by p/m 1 and then by 3 let m know what you get ? what would be the possible zeros ?

OpenStudy (steve816):

1, 2, 4, 8, 16, 32, 64, 1/3, 2/3, 4/3, 8/3, 16/3, 32/3, 64/3

OpenStudy (steve816):

What do i do from here, please help!

Nnesha (nnesha):

test each number remember all of them are plus/minus so start with p/m 1 \[\rm f(\color{ReD}{-1})=3x^2+4x^3-36x^2+64\]\[\rm f(\color{ReD}{1})=3x^2+4x^3-36x^2+64\]\[\rm f(\color{ReD}{-2})=3x^2+4x^3-36x^2+64\]\[\rm f(\color{ReD}{2})=3x^2+4x^3-36x^2+64\] or do synthetic division

OpenStudy (steve816):

NEGATIVES TOO? That is going to take forever :( I think i am going to give up on this problem.

Nnesha (nnesha):

i don't which one is easy for you  if you get 0 remainder then that would be the real root